# out-of-domain OC20 dataset Open Catalyst 2020 dataset

### v1

@@ Methods
The amount of data used to train the models directly impacts the cost to train and to iterate between architectures and ablations. Gasteiger and colleagues have shown the OC20-2M subset to be representative of the full OC20 dataset, primarily because it preserved the underlying chemical diversityGasteiger et al. 2022. Other approaches use more complex stratified sampling (based on feature-space clustering) to ensure that diverse, high-energy, and uncommon configurations are explicitly captured to improve model robustnessQi et al. 2023. Given the precalculated training data, we explored the opportunity to reduce model training costs by sampling the frames along adsorbate-slab relaxation trajectories. Sampling was performed using force-stratified selections of the trajectory to obtain a representative distribution of systems. For consistency, models were trained for a nearly constant number of total gradient updates, approximately equivalent to the number of steps in one epoch of the largest split. Further, the data ablation for sampling frames from trajectories (Figure 4a) was designed to probe redundancy in highly autocorrelated data and thus only applied to the relaxations and MD data; the rattled and transition state data were included entirely for each model. To enable a cleaner evaluation of the data cost-benefit trade-off for tuning, without confounding the results with the model’s ability to learn new, unseen elements, we restricted the subsampling experiments to AQCat25 systems with elements already seen by the 31M pretrained model. Figure 4a reports the validation MAE from the final training checkpoint, which highlights the risk of overfitting. Conversely, Figure 4b plots the MAE from the best-performing model during training for each slab count (averaged over all k values) against the data generation cost. As can be seen in Figure 4a, for small numbers of slabs, sampling a subset of frames rather than using the full trajectory actually leads to better force MAE values. This is likely due to a high propensity for overfitting, which is shown in Figure 10 and remedied by sampling. At larger slab numbers, the differences are minor but the cost to train will be lower if sampling is performed. For energy MAE there is not a substantial trend with changing sample size (see Figure 11). Therefore, sampling frames is a useful cost-saving strategy. We adopted a subsampling approach for the model adaptation experiments presented in subsequent sections. However, for those experiments, we employed random sampling per trajectory rather than the force-stratified approach. We found this yielded slightly improved performance, likely due to increasing the representation of low-force, near-equilibrium frames that are critical for downstream adsorption energy tasks.
@@
These patterns follow from standard behavior under distribution shift and multi-domain supervision. OC20 is broader and uses different DFT settings than AQCat25. Fine-tuning only on AQCat25 moves the parameters toward that narrower distribution and forgets OC20-specific features. Adding replay during cotuning without FiLM counteracts forgetting but also pulls the solution toward OC20 conventions, which explains the performance degradation on AQCat25 forces in Figure 6c. Introducing FiLM provides a framework to distinguish these distributions, which rectifies the decrease in the force metric. Starting from scratch changes the optimization path. The performance trends are strongly dependent on the OC20 data size used. Unsuprisingly, at 0M, jointly tuning clearly outperforms jointly training from scratch, but as OC20 data is added, their performance becomes sensitive to the metric. While tuning holds a slight advantage at a very strict 0.1 eV low-error threshold, the models cotrained from scratch show an advantage at higher cutoffs (e.g., 0.6-0.8 eV), indicating they are more effective at capturing outliers (see Figures 16a-f. FiLM makes the domain information explicit. Conditioning on spin and fidelity yields feature rescaling that reduces gradient interference between magnetic vs. non-magnetic and high- vs. low-fidelity cases.
@@
We also explored the robustness and generalizability of models to form a more complete assessment of model usability. To do this, we constructed an additional validation set aimed at assessing the ability of the models to identify the global minimum energy for a given adsorbate-slab combination in line with the approach presented by Lan et alLan et al. 2023. We further explored differences in model performance when segmenting the data by interesting splits, namely the material type (metal-only, non-metal, metalloid, and metalloid+non-metal), whether spin was on or off, and whether the elements in a material were all included in OC20 or not.
@@
The ultimate use of the MLIPs trained here will be for practical catalyst discovery where an important figure of merit is the global minimum adsorption energy. We explored this using the dense DFT validation set with 50 relaxations each for 109 adsorbate-slab combinations. We performed ML relaxation inference starting from the same initial configurations as DFT. The relaxed states were filtered using the same algorithms presented by Lan et al.Lan et al. 2023 to find desorption, dissociation, intercalation, and significant surface change. Figure 7 compares the performance of three 31M parameter models on this task, using only the ML-predicted energies without DFT single-points on the ML-relaxed structures. On the left is the pretrained 31M EV2 model (trained on OC20 All+MD), taken from the publicly available fairchem checkpoint. For this model, inference on systems containing new elements were omitted since performance would be poor. In the center is a directly fine-tuned EV2 model. On the right is the EV2-in+midFilM model, which was cotrained using 20M OC20 examples and the AQCat25 dataset. As a point of comparison, when the OC20denseLan et al. 2023 dataset was released, the EV2 model was not available. The best performing model was eSCN-MD-Large and on this task it had a 56.5% success rate with an energy MAE of 0.17 eVLan et al. 2023. Here, success rate is defined as the percent of systems where the minimum adsorption energy found by ML is within 0.1 eV of the DFT value. ML success metrics alone were included in a later release as 60.8% for an EV2 model of unspecified size, 68.4% for the UMA-S model, 71.1% for the UMA-M model, and 74.4% for the UMA-LWood et al. 2025. The success metric and MAEs have been annotated on the plots. We see the trend we would expect to see between models, with increasing performance from left to right. This further validates the usefulness of the models, and also supports the fact that the loss function and training metrics are well designed and correlate with our downstream use case.
@@ Conclusion
O.A.: Model implementation, experimental design, training, ablations, idea conceptualization, data processing, data analysis, data visualization, writing, editing. B.W.: Dataset generation, idea conceptualization, data analysis, data visualization, writing, editing. A.R.S.: Project leadership, idea conceptualization, writing, editing.
@@ Acknowledgements
The authors acknowledge SandboxAQ leadership, especially Ang Xiao, Adam Lewis, Arman Zaribafiyan, Jeff Graf, Nadia Harhen, Andrew McLaughlin, and Jack Hidary, for their support of this research. The authors thank Joseph Gauthier, Kevin Ryczko, Tom Ludwig, Jia-Min Chu, Tyler Sours, Jiyoon Kim, and Jens Nørskov for valuable discussions, code reviews, and feedback. The authors recognize computational and engineering support from SungYeon Kim, Rudi Plesch, and the Nvidia DGX team.
@@
Figure 9 shows a comparison between the frequency with which elements occur in the AQCat25 and OC20 datasets. There are some notable differences like the presence of the six additional elements in AQCat25, the higher relative presense of boron in AQCat25, and the lower presense of Tc in AQCat25.
@@
Figure 11 shows: (a) Energy MAE which does not show a substantial trend with changing k for random subsampling, but is improved by increasing the number of slabs. (b) The AQCat25 validation force MAE for the pretrained 31M parameter Equiformer v2 model across stratified force bins, which shows roughly the same trend as energy: performance decreases on higher force systems with the exception of very high force frames which have better performance. (c) The AQCat25 validation force MAE for naively fintuned models using different values of first k samples of the AQCat25 dataset to fine-tune.
@@
Table 7: Architectural hyperparameters for the EquiformerV2 models, including FiLM specifics. We refer to the EquiformerV2 paper Liao et al. 2023 for a complete description of all architectural components, including normalization and activation functions.
@@
The results looking at the dense dataset but split over whether the system was spin off (Figure 14 or on and whether the system contains new elements reveals that there are not any strong discrepancies between these groups. This is in alignment with Figure 8.
@@
Figure 17: Performance of contuned and cotrained models on test and val for energies (top) and forces (bottom).

### v2

@@ Methods
The amount of data used to train the models directly impacts the cost to train and to iterate between architectures and ablations. Gasteiger and colleagues have shown the OC20-2M subset to be representative of the full OC20 dataset, primarily because it preserved the underlying chemical diversitygasteiger2022gemnetoc. Other approaches use more complex stratified sampling (based on feature-space clustering) to ensure that diverse, high-energy, and uncommon configurations are explicitly captured to improve model robustnessqi2023direct. Given the precalculated training data, we explored the opportunity to reduce model training costs by sampling the frames along adsorbate-slab relaxation trajectories. Sampling was performed using force-stratified selections of the trajectory to obtain a representative distribution of systems. For consistency, models were trained for a nearly constant number of total gradient updates, approximately equivalent to the number of steps in one epoch of the largest split. Further, the data ablation for sampling frames from trajectories (Figure 4a) was designed to probe redundancy in highly autocorrelated data and thus only applied to the relaxations and MD data; the rattled and transition state data were included entirely for each model. To enable a cleaner evaluation of the data cost-benefit trade-off for tuning, without confounding the results with the model’s ability to learn new, unseen elements, we restricted the subsampling experiments to AQCat25 systems with elements already seen by the 31M pretrained model. Figure 4a reports the validation MAE from the final training checkpoint, which highlights the risk of overfitting. Conversely, Figure 4b plots the MAE from the best-performing model during training for each slab count (averaged over all k values) against the data generation cost. As can be seen in Figure 4a, for small numbers of slabs, sampling a subset of frames rather than using the full trajectory actually leads to better force MAE values. This is likely due to a high propensity for overfitting, which is shown in Figure 10 and remedied by sampling. At larger slab numbers, the differences are minor but the cost to train will be lower if sampling is performed. For energy MAE there is not a substantial trend with changing sample size (see Figure 11). Therefore, sampling frames is a useful cost-saving strategy. We adopted a subsampling approach for the model adaptation experiments presented in subsequent sections. However, for those experiments, we employed random sampling per trajectory rather than the force-stratified approach. We found this yielded slightly improved performance, likely due to increasing the representation of low-force, near-equilibrium frames that are critical for downstream adsorption energy tasks.
@@
These patterns follow from standard behavior under distribution shift and multi-domain supervision. OC20 is broader and uses different DFT settings than AQCat25. Fine-tuning only on AQCat25 moves the parameters toward that narrower distribution and forgets OC20-specific features. Adding replay during cotuning without FiLM counteracts forgetting but also pulls the solution toward OC20 conventions, which explains the performance degradation on AQCat25 forces in Figure 6c. Introducing FiLM provides a framework to distinguish these distributions, which rectifies the decrease in the force metric. Starting from scratch changes the optimization path. The performance trends are strongly dependent on the OC20 data size used. Unsurprisingly, at 0M, jointly tuning clearly outperforms jointly training from scratch, but as OC20 data is added, their performance becomes sensitive to the metric. While tuning holds a slight advantage at a very strict 0.1 eV low-error threshold, the models cotrained from scratch show an advantage at higher cutoffs (e.g., 0.6-0.8 eV), indicating they are more effective at capturing outliers (see Figures 16a-f. FiLM makes the domain information explicit. Conditioning on spin and fidelity yields feature rescaling that reduces gradient interference between magnetic vs. non-magnetic and high- vs. low-fidelity cases.
@@
We also explored the robustness and generalizability of models to form a more complete assessment of model usability. To do this, we constructed an additional validation set aimed at assessing the ability of the models to identify the global minimum energy for a given adsorbate-slab combination in line with the approach presented by Lan et allan2023adsorbml. We further explored differences in model performance when segmenting the data by interesting splits, namely the material type (metal-only, non-metal, metalloid, and metalloid+non-metal), whether spin was on or off, and whether the elements in a material were all included in OC20 or not.
@@
The ultimate use of the MLIPs trained here will be for practical catalyst discovery where an important figure of merit is the global minimum adsorption energy. We explored this using the dense DFT validation set with 50 relaxations each for 109 adsorbate-slab combinations. We performed ML relaxation inference starting from the same initial configurations as DFT. The relaxed states were filtered using the same algorithms presented by Lan et al.lan2023adsorbml to find desorption, dissociation, intercalation, and significant surface change. Figure 7 compares the performance of three 31M parameter models on this task, using only the ML-predicted energies without DFT single-points on the ML-relaxed structures. On the left is the pretrained 31M EV2 model (trained on OC20 All+MD), taken from the publicly available fairchem checkpoint. For this model, inference on systems containing new elements were omitted since performance would be poor. In the center is a directly fine-tuned EV2 model. On the right is the EV2-in+midFilM model, which was cotrained using 20M OC20 examples and the AQCat25 dataset. As a point of comparison, when the OC20denselan2023adsorbml dataset was released, the EV2 model was not available. The best performing model was eSCN-MD-Large and on this task it had a 56.5% success rate with an energy MAE of 0.17 eVlan2023adsorbml. Here, success rate is defined as the percent of systems where the minimum adsorption energy found by ML is within 0.1 eV of the DFT value. ML success metrics alone were included in a later release as 60.8% for an EV2 model of unspecified size, 68.4% for the UMA-S model, 71.1% for the UMA-M model, and 74.4% for the UMA-Lwood2025family. The success metric and MAEs have been annotated on the plots. We see the trend we would expect to see between models, with increasing performance from left to right. This further validates the usefulness of the models, and also supports the fact that the loss function and training metrics are well designed and correlate with our downstream use case.
@@ Conclusion
O.A.: Model implementation, experimental design, training, ablations, idea conceptualization, data processing, data analysis, data visualization, writing, editing. B.W.: Dataset generation, idea conceptualization, data analysis, data visualization, writing, editing. A.R.S.: Project leadership, idea conceptualization, writing, editing. S.K., R.P., P.A., A.N.: Computing infrastructure. T.S., J.C., T.L., R.W., S.A., A.R., A.F., C.W., A.W., T.M., K.R.: Code validation, manuscript review.
@@ Acknowledgements
The authors acknowledge SandboxAQ leadership, especially Ang Xiao, Adam Lewis, Arman Zaribafiyan, Jeff Graf, Takeshi Yamazaki, Nadia Harhen, Andrew McLaughlin, and Jack Hidary, for their support of this research. The authors thank Joseph Gauthier and Jens Nørskov for valuable discussions and feedback. The authors recognize computational and engineering support from the Nvidia DGX team.
@@
Figure 9 shows a comparison between the frequency with which elements occur in the AQCat25 and OC20 datasets. There are some notable differences like the presence of the six additional elements in AQCat25, the higher relative presence of boron in AQCat25, and the lower presence of Tc in AQCat25.
@@
Figure 11 shows: (a) Energy MAE which does not show a substantial trend with changing k for random subsampling, but is improved by increasing the number of slabs. (b) The AQCat25 validation force MAE for the pretrained 31M parameter Equiformer v2 model across stratified force bins, which shows roughly the same trend as energy: performance decreases on higher force systems with the exception of very high force frames which have better performance. (c) The AQCat25 validation force MAE for naively finetuned models using different values of first k samples of the AQCat25 dataset to fine-tune.
@@
Table 7: Architectural hyperparameters for the EquiformerV2 models, including FiLM specifics. We refer to the EquiformerV2 paper liao2023equiformerv2 for a complete description of all architectural components, including normalization and activation functions.
@@
Figure 14 splits the dense dataset results over whether the system was spin off or on and whether the system contains new elements, revealing that there are not any strong discrepancies between these groups. This is in alignment with Figure 8.
@@
Figure 17: Performance of cotuned and cotrained models on test and val for energies (top) and forces (bottom).